OpenStudy (rissyroo118):

Question 1.1. Factor 2x2 + 8x + 8 medal given x2+ 4x + 4 2(x + 2) 2(x + 2)2 (x + 2)

OpenStudy (acxbox22):

first factor a two out

OpenStudy (rissyroo118):

2x+9x+8=11x+8

OpenStudy (rissyroo118):

@mathmate

OpenStudy (mathmate):

Factor out the common of 2 and use perfect square. \(2x2 + 8x + 8 = 2(x^2+4x+4)\) perfect square: \((a+b)^2=a^2+2ab+b^2\) where a=x, b=2